If 50 randomly selected college students take the examination, what is the probability that the mean time it takes the group to complete the test will be less than 43 minutes?

1
Expert's answer
2021-11-17T15:47:43-0500

"P(\bar X<43)=P(Z<\frac{43-46.2}{\frac{8}{\sqrt{50}}})=P(Z<-2.83)=0.0023."